# Umlaut information

### v1

@@ I Introduction
$U(X;Y)$ $=-H(P_{X})-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log P_{XY}(x,y)\right).$ (6)
This formulation allows us to prove the additivity of the umlaut information under tensor product. In turn, this then implies an operational interpretation of the umlaut information in composite hypothesis testing. More precisely, for a joint probability distribution $P_{XY}$ over $\mathcal{X}\times\mathcal{Y}$ with marginal $P_{X}$, we establish that the umlaut information $U(X;Y)$ governs the exponential decaying rate of the type I-error while maintaining a type II-error smaller than a constant $\varepsilon>0$ for the problem of testing
$H_{0}:R_{X^{n}Y^{n}}=P_{XY}^{\times n}$ versus $H_{1}:R_{X^{n}Y^{n}}=P_{X}^{\times n}Q_{Y^{n}}$, (7)
where $P_{Y^{n}}$ is an arbitrary probability distribution on $\mathcal{Y^{n}}$.
@@
$U(\pazocal{W}_{1}\times\pazocal{W}_{2})=U(\pazocal{W}_{1})+U(\pazocal{W}_{2}).$ (9)
@@
$\varepsilon(\exp(rn),\pazocal{W}^{\times n})\simeq\exp(-nE(r,\pazocal{W})+o(n)),$ (10)
where $E(r,\pazocal{W})$ is known as the error exponent (sometimes also termed reliability function). Although the standard setting of communication relies on encoders and decoders without additional assistance, many converse results can be proved for more general, assisted coding schemes. In particular, Polyanskiy, Poor and Verdú [7] introduced the meta-converse bound that implies many well-known converse results in the literature, and [8] later showed that this bound in fact exactly corresponds to the setting of coding assisted by non-signalling (NS) correlations. Now, in the zero-rate regime where $\frac{1}{n}\log M\rightarrow 0$, we establish that the non-signalling–assisted error exponent is quantified by the channel umlaut information. That is, we find
$E^{\rm NS}(0^{+},\pazocal{W})=E_{\mathrm{sp}}(0^{+},\pazocal{W})=U(\pazocal{W}),$ (11)
@@
Returning to the unassisted setting, we can relax the decoding requirement to the list decoding where the error occurs only if the original message does not belong to the list of messages (of size $L\geq 1$) returned by the decoder. In this setting, the zero-rate unassisted reliability function is [9, Theorem] (see also [10] for further discussions)
$E^{\emptyset}_{L}(0^{+},\pazocal{W})=\max_{P_{X}}\sum_{x_{1},\dots,x_{L+1}}P_{X}(x_{1})\cdots P_{X}(x_{L+1})\left(-\log\sum_{y\in\mathcal{Y}}\sqrt[L+1]{\pazocal{W}(y|x_{1})\cdots\pazocal{W}(y|x_{L+1})}\right).$ (12)
As the second operational interpretation of the channel lautum information, we show that for all $L\geq 1$, the channel umlaut information $U(\pazocal{W})$ provides an upper bound to the unassisted zero-rate error exponent $E_{L}(0^{+},\pazocal{W})$ and establish that the rate $U(\pazocal{W})$ is in fact achieved in the large list limit $L\to\infty$.
@@ II Notation and Preliminaries
$D(P\|Q)\coloneqq\int_{\mathcal{X}}\,\log\frac{\mathrm{d}P}{\mathrm{d}Q}\,\mathrm{d}P$ (13)
if $P\ll Q$ and +$\infty$+ ∞ otherwise. In the former case the Radon-Nikodym derivative $\frac{\mathrm{d}P}{\mathrm{d}Q}$ is well-defined $P$-everywhere. We will mostly work with discrete spaces $\mathcal{X}$ with $\Sigma_{\mathcal{X}}$ the power set of $\mathcal{X}$, in which case the relative entropy takes the form
$D(P\|Q)=\sum_{x\in\mathcal{X}}P(x)\log\frac{P(x)}{Q(x)}$ (14)
and $P$, $Q$ are the probability mass functions defined on $\mathcal{X}$. We use standard notions as summarised in Table 1 to describe random variables in this case.
@@ III.1 Definition and basic properties
Despite their superficial similarity, there are many differences between mutual and lautum information: for instance, $I(X\!:\!Y)$ can be expressed as a linear combination of entropic measure [12], while the same property is not true for $L(X\!:\!Y)$ [3]. Furthermore, it is well known that the mutual information can be written in the following variational forms:
$I(X\!:Y\!)=D(P_{XY}\|P_{X}P_{Y})=\min_{Q_{Y}}D(P_{XY}\|P_{X}Q_{Y})=\min_{Q_{X}}\min_{Q_{Y}}D(P_{XY}\|Q_{X}Q_{Y}).$ (20)
@@
$P_{XY}(x,y)=\begin{cases}0&x=y=0\\ \frac{1}{3}&\text{otherwise,}\end{cases}$ (21)
@@
    Data-processing inequality: Consider a Markov chain $X-Y-Z$. Then,
@@
For (3), it suffices to show the first inequality since $X-Y-Z$ implies $Z-Y-X$. The Markov condition than implies the existence of a channel $\pazocal{W}_{Z|Y}$ such that $P_{XZ}(x,z)=\sum_{y\in\mathcal{Y}}P_{XY}(x,y)\pazocal{W}(z|y)$. Using the monotonicity of relative entropy under channels we arrive at
@@
Definition 2 can be generalised in terms of the Rényi $\alpha$-relative entropies. These quantities previously appeared in [6, Appendix E] as a tool used to obtain bounds on the error probability of (quantum) state exclusion.
Definition 4 ((Rényi $\alpha$-umlaut information)).
$U_{\alpha}(X;Y)$ $\coloneqq\min_{Q_{Y}}D_{\alpha}(P_{X}Q_{Y}\|P_{XY})$ (25)
$D_{\alpha}(P\|Q)\coloneqq\frac{1}{\alpha-1}\log\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)Q_{X}^{1-\alpha}(x)\,.$ (26)
Lemma 5.
Given two random variables $X$ and $Y$, it holds that
$\limsup_{\alpha\to 1^{-}}U_{\alpha}(X;Y)=U(X;Y).$ (27)
Proof.
Let $P_{XY}$ be the joint probability distribution of $X$ and $Y$. Since $\alpha\mapsto D_{\alpha}(p\|q)$ is a monotonically increasing function, we can write
$\limsup_{\alpha\to 1^{-}}U_{\alpha}(X;Y)=\sup_{\alpha<1}\min_{Q_{Y}}D_{\alpha}(P_{X}Q_{Y}\|P_{XY})$ (28)
By the Mosonyi–Hiai minimax theorem [13, Corollary A2], we can rewrite
$\limsup_{\alpha\to 1^{-}}U_{\alpha}(X;Y)=\min_{Q_{Y}}\sup_{\alpha<1}D_{\alpha}(P_{X}Q_{Y}\|P_{XY})\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}\min_{Q_{Y}}D(P_{X}Q_{Y}\|P_{XY})=U(X;Y),$ (29)
where in (i) we have used that the Rényi relative entropies converge to the Kullback–Leibler divergence as $\alpha\to 1^{-}$. ∎
@@ III.2 Closed form and optimal marginal
In this subsection we find the minimiser appearing in our variational definition of the umlaut information. As a consequence, we can write down a closed formula for $U(X;Y)$ and prove additivity. The same computations can be done in the case of the Rényi $\alpha$-umlaut information: also in this case additivity holds.
@@
$\ddot{P}_{Y}(y)\coloneqq\frac{1}{Z[P_{XY}]}\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log P_{XY}(x,y)\right)\qquad\forall y\in\mathcal{Y},$ (30)
@@
An alternative argument to derive the closed form of the umlaut information presented in the above Theorem 7 can be obtained by carefully taking the limit $\alpha\to 1$ in [6, Lemma 35]. Using the explicit formula for $U$, it is elementary to prove its additivity.
@@
$\ddot{P}_{YY^{\prime}}(y,y^{\prime})$ $=\frac{1}{Z[P_{XY}P^{\prime}_{X^{\prime}Y^{\prime}}]}\exp\left(\sum_{\begin{subarray}{c}x\in\mathcal{X}\\ x^{\prime}\in\mathcal{X^{\prime}}\end{subarray}}P_{X}(x)P_{X^{\prime}}(x^{\prime})\log\left(P_{XY}(x,y)P^{\prime}_{X^{\prime}Y^{\prime}}(x^{\prime},y^{\prime})\right)\right)$ (35)
@@
$U(XX^{\prime};YY^{\prime})$ $=D(P_{X}P_{X^{\prime}}\ddot{P}_{YY^{\prime}}\|P_{XY}P^{\prime}_{X^{\prime}Y^{\prime}})$ (36)
@@
For any $\alpha\in(0,1)$, given a joint probability distribution $P_{XY}\in\mathcal{P}(\mathcal{X}\times\mathcal{Y})$, it holds that
$U_{\alpha}(X;Y)=D_{\alpha}\big{(}P_{X}\ddot{P}^{(\alpha)}_{Y}\,\big{\|}\,P_{XY}\big{)}=-\log\sum_{y\in\mathcal{Y}}P_{Y}(y)\left(\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{X|Y}^{1-\alpha}(x|y)\right)^{\frac{1}{1-\alpha}},$ (37)
where $P_{X}$ is the marginal of $P_{XY}$ on $\mathcal{X}$ and
$\ddot{P}_{Y}^{(\alpha)}(y)\coloneqq\frac{1}{Z_{\alpha}[P_{XY}]}\left(\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{XY}^{1-\alpha}(x,y)\right)^{\frac{1}{1-\alpha}},$ (38)
where $Z_{\alpha}[P_{XY}]$ is the normalisation constant making $\ddot{P}_{Y}^{(\alpha)}$ a probability distribution on $\mathcal{Y}$. In particular, $U_{\alpha}$ is additive.
@@
Let us rewrite
$U_{\alpha}(X;Y)$ $=\min_{Q_{Y}}\frac{1}{\alpha-1}\log\sum_{\begin{subarray}{c}x\in\mathcal{X}\\ y\in\mathcal{Y}\end{subarray}}P_{X}^{\alpha}(x)Q_{Y}^{\alpha}(y)P_{XY}^{1-\alpha}(x,y)$ (39)
@@
$U_{\alpha}(X;Y)$ $=\frac{1}{\alpha-1}\log\max_{\begin{subarray}{c}f\geq 0,\\ \|f\|_{1/\alpha}=1\end{subarray}}\langle f,g\rangle=\frac{1}{\alpha-1}\log\|g\|_{\frac{1}{1-\alpha}}=-\log\left\|\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{XY}^{1-\alpha}(x,\,\cdot\,)\right\|_{\frac{1}{1-\alpha}}^{\frac{1}{1-\alpha}},$ (41)
@@
$f=\frac{\left(\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{XY}^{1-\alpha}(x,\,\cdot\,)\right)^{\frac{\alpha}{1-\alpha}}}{\left[\sum_{y\in\mathcal{Y}}\left(\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{XY}^{1-\alpha}(x,y)\right)^{\frac{1}{1-\alpha}}\right]^{\alpha}}\eqcolon\left(\ddot{P}_{Y}^{(\alpha)}\right)^{\alpha}.$ (42)

### v2

@@ I Introduction
$\hskip 0.0ptU(X;Y)$ $=-H(P_{X})-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log P_{XY}(x,y)\right).$ (6)
This formulation allows us to prove the additivity of the umlaut information under tensor product. In turn, this then implies an operational interpretation of the umlaut information in composite hypothesis testing. More precisely, for a joint probability distribution $P_{XY}$ over $\mathcal{X}\times\mathcal{Y}$ with marginal $P_{X}$, we establish that the umlaut information $U(X;Y)$ governs the exponential decay rate of the type II error probability while maintaining a type I error probability smaller than a constant $\varepsilon>0$ for the problem of testing
$H_{0}:R_{X^{n}Y^{n}}=P_{X}^{\times n}Q_{Y^{n}}$ versus $H_{1}:R_{X^{n}Y^{n}}=P_{XY}^{\times n}$, (7)
where $Q_{Y^{n}}$ is an arbitrary probability distribution on $\mathcal{Y^{n}}$.
@@
$\hskip 0.0ptU(\pazocal{W}_{1}\times\pazocal{W}_{2})=U(\pazocal{W}_{1})+U(\pazocal{W}_{2}).$ (9)
A similar property holds also for variants of the umlaut information defined in terms of Rényi $\alpha$-divergences, which we also introduce and study in detail.
@@
$\hskip 0.0pt\varepsilon(\exp(rn),\pazocal{W}^{\times n})\simeq\exp(-nE(r,\pazocal{W})+o(n)),$ (10)
where $E(r,\pazocal{W})$ is known as the error exponent (sometimes also termed reliability function). Although the standard setting of communication relies on encoders and decoders without additional assistance, many converse results can be proved for more general, assisted coding schemes. In particular, Polyanskiy, Poor and Verdú [7] introduced the meta-converse bound that implies many well-known converse results in the literature, and Matthews [8] later showed that this bound in fact exactly corresponds to the setting of coding assisted by non-signalling (NS) correlations. Now, in the zero-rate regime where $\frac{1}{n}\log M\rightarrow 0$, we establish that the non-signalling–assisted error exponent is quantified by the channel umlaut information. That is, we find
$\hskip 0.0ptE^{\rm NS}(0^{+},\pazocal{W})=E_{\mathrm{sp}}(0^{+},\pazocal{W})=U(\pazocal{W}),$ (11)
@@
Returning to the unassisted setting, we can relax the decoding requirement to the list decoding where the error occurs only if the original message does not belong to the list of messages (of size $L\geq 1$) returned by the decoder. In this setting, the zero-rate unassisted reliability function is [9] (see also [10] for further discussions)
[…]
As the second operational interpretation of the channel lautum information, we show that for all $L\geq 1$, the channel umlaut information $U(\pazocal{W})$ provides an upper bound to the unassisted zero-rate error exponent $E_{L}(0^{+},\pazocal{W})$ and establish that the rate $U(\pazocal{W})$ is in fact achieved in the large list limit $L\to\infty$. This is similar to, but different from, prior literature that studied an asymptotic limit of exponentially large list sizes and also connected it with the sphere packing bound [11].
@@ II Notation and Preliminaries
$\hskip 0.0ptD(P\|Q)\coloneqq\int_{\mathcal{X}}\,\log\frac{\mathrm{d}P}{\mathrm{d}Q}\,\mathrm{d}P$ (13)
if $P\ll Q$ and $\infty$ otherwise. In the former case the Radon-Nikodym derivative $\frac{\mathrm{d}P}{\mathrm{d}Q}$ is well-defined $P$-everywhere. We will mostly work with finite spaces $\mathcal{X}$, with $\Sigma_{\mathcal{X}}$ the power set of $\mathcal{X}$, in which case the relative entropy takes the form
$D(P\|Q)=\sum_{x\in\mathcal{X}}P(x)\log\frac{P(x)}{Q(x)}\,$, (14)
and $P$, $Q$ are the probability mass functions defined on $\mathcal{X}$. The discussion of the case of a continuous probability density function is relegated to Appendix B. We use standard notation to describe random variables, as summarised in Table 1.
@@ III.1 Definition and basic properties
Despite their superficial similarity, there are many differences between mutual and lautum information: for instance, I(X:Y)I(X\!:\!Y) can be expressed as a linear combination of entropic measures [13], while the same property is not true for L(X:Y)L(X\!:\!Y) [3]. Furthermore, it is well known that the mutual information can be written in the following variational forms:
$\hskip 0.0ptI(X\!:Y\!)=D(P_{XY}\|P_{X}P_{Y})=\min_{Q_{Y}}D(P_{XY}\|P_{X}Q_{Y})=\min_{Q_{X}}\min_{Q_{Y}}D(P_{XY}\|Q_{X}Q_{Y}).$ (20)
@@
[…]
@@
    Data-processing inequality: Consider a Markov chain XYZX-Y-Z. Then,
@@
For (3), it suffices to show the first inequality since XYZX-Y-Z implies ZYXZ-Y-X. The Markov condition than implies the existence of a channel $\pazocal{W}_{Z|Y}$ such that $P_{XZ}(x,z)=\sum_{y\in\mathcal{Y}}P_{XY}(x,y)\pazocal{W}(z|y)$. Using the monotonicity of relative entropy under channels we arrive at
@@ III.2 Closed form and optimal marginal
In this subsection we find the minimiser appearing in our variational definition of the umlaut information. As a consequence, we can write down a closed formula for $U(X;Y)$ and prove additivity. The same computations can be done in the case of the Rényi $\alpha$-umlaut information; also in this case additivity holds.
@@
$\hskip 0.0pt\ddot{P}_{Y}(y)\coloneqq\frac{1}{Z[P_{XY}]}\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log P_{XY}(x,y)\right)\qquad\forall y\in\mathcal{Y},$ (25)
@@
An alternative argument to derive the closed form of the umlaut information presented in the above Proposition 5 can be obtained by carefully taking the limit $\alpha\to 1$ in [6, Lemma 38]. Using the explicit formula for $U$, it is elementary to prove its additivity.
@@
$\hskip 0.0pt\ddot{P}_{YY^{\prime}}(y,y^{\prime})$ $=\frac{1}{Z[P_{XY}P^{\prime}_{X^{\prime}Y^{\prime}}]}\exp\left(\sum_{\begin{subarray}{c}x\in\mathcal{X}\\ x^{\prime}\in\mathcal{X^{\prime}}\end{subarray}}P_{X}(x)P_{X^{\prime}}(x^{\prime})\log\left(P_{XY}(x,y)P^{\prime}_{X^{\prime}Y^{\prime}}(x^{\prime},y^{\prime})\right)\right)$ (30)
@@
$\hskip 0.0ptU(XX^{\prime};YY^{\prime})$ $=D(P_{X}P_{X^{\prime}}\ddot{P}_{YY^{\prime}}\|P_{XY}P^{\prime}_{X^{\prime}Y^{\prime}})$ (31)
@@
Remark 7.
We can manipulate the closed-form expression for the umlaut information in order to find an alternative form:
$\hskip 0.0ptU(X;Y)$ $=-H(P_{X})-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log P_{XY}(x,y)\right)$ (32)
$=-\log\sum_{y\in\mathcal{Y}}P_{Y}(y)\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log\frac{P_{XY}(x,y)}{P_{X}(x)P_{Y}(y)}\right)$
$=-\log\mathbb{E}_{Z\sim P_{Y}}\left[\exp\left(-D(P_{X}\|P_{X|Y=Z})\right)\right].$
III.3 Rényi divergence variant
The definition the umlaut information (Definition 2) can be generalised in terms of the Rényi $\alpha$-relative entropies. These quantities previously appeared in [6, Appendix E] as a tool used to obtain bounds on the error probability of (quantum) state exclusion.
Definition 8 ((Rényi $\alpha$-umlaut information)).
$\hskip 0.0ptU_{\alpha}(X;Y)$ $\coloneqq\min_{Q_{Y}}D_{\alpha}(P_{X}Q_{Y}\|P_{XY})$ (33)
$\hskip 0.0ptD_{\alpha}(P\|Q)\coloneqq\frac{1}{\alpha-1}\log\sum_{x\in\mathcal{X}}P^{\alpha}(x)Q^{1-\alpha}(x)\,.$ (34)
Lemma 9.
Given two random variables $X$ and $Y$, it holds that
$\hskip 0.0pt\lim_{\alpha\to 1^{-}}U_{\alpha}(X;Y)=U(X;Y)$ (35)
and
$\hskip 0.0pt\lim_{\alpha\to 1^{+}}U_{\alpha}(X;Y)=U(X;Y).$ (36)
Proof.
Let $P_{XY}$ be the joint probability distribution of $X$ and $Y$. Since $\alpha\mapsto D_{\alpha}(p\|q)$ is a monotonically increasing function, for $0<\alpha<1$ we can write
$\hskip 0.0pt\lim_{\alpha\to 1^{-}}U_{\alpha}(X;Y)=\sup_{\alpha<1}\min_{Q_{Y}}D_{\alpha}(P_{X}Q_{Y}\|P_{XY})$ (37)
By the Mosonyi–Hiai minimax theorem [14, Corollary A2], we can rewrite
$\hskip 0.0pt\lim_{\alpha\to 1^{-}}U_{\alpha}(X;Y)=\min_{Q_{Y}}\sup_{\alpha<1}D_{\alpha}(P_{X}Q_{Y}\|P_{XY})\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}\min_{Q_{Y}}D(P_{X}Q_{Y}\|P_{XY})=U(X;Y),$ (38)
where in (i) we have used that the Rényi relative entropies converge to the Kullback–Leibler divergence as $\alpha\to 1^{-}$. For $\alpha>1$, leveraging again on the monotonicity of $\alpha\mapsto D_{\alpha}(p\|q)$, we can easily compute
$\hskip 0.0pt\lim_{\alpha\to 1^{+}}U_{\alpha}(X;Y)$ $=\inf_{\alpha<1}\min_{Q_{Y}}D_{\alpha}(P_{X}Q_{Y}\|P_{XY})=\min_{Q_{Y}}\inf_{\alpha<1}D_{\alpha}(P_{X}Q_{Y}\|P_{XY})$ (39)
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(iii)}}}}{{=}}\min_{Q_{Y}}D(P_{X}Q_{Y}\|P_{XY})=U(X;Y),$
where in (iii) we have used that the Rényi relative entropies converge to the Kullback–Leibler divergence as $\alpha\to 1^{+}$. ∎
@@ III.3 Rényi divergence variant
For any $\alpha\in(0,1)\cup(1,\infty)$, given a joint probability distribution $P_{XY}\in\mathcal{P}(\mathcal{X}\times\mathcal{Y})$, it holds that
$\hskip 0.0ptU_{\alpha}(X;Y)=D_{\alpha}\big{(}P_{X}\ddot{P}^{(\alpha)}_{Y}\,\big{\|}\,P_{XY}\big{)}=-\log\sum_{y\in\mathcal{Y}}P_{Y}(y)\left(\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{X|Y}^{1-\alpha}(x|y)\right)^{\frac{1}{1-\alpha}},$ (40)
where $P_{X}$ is the marginal of $P_{XY}$ on $\mathcal{X}$ and, assuming $U_{\alpha}(X;Y)<+\infty$,
$\hskip 0.0pt\ddot{P}_{Y}^{(\alpha)}(y)\coloneqq\frac{1}{Z_{\alpha}[P_{XY}]}\left(\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{XY}^{1-\alpha}(x,y)\right)^{\frac{1}{1-\alpha}},$ (41)
where $Z_{\alpha}[P_{XY}]$ is the normalisation constant making $\ddot{P}_{Y}^{(\alpha)}$ a probability distribution on $\mathcal{Y}$, with the conventions $0^{-1}=+\infty$ and $\infty^{-1}=0$. If $U_{\alpha}(X;Y)=+\infty$, then $\ddot{P}_{Y}^{(\alpha)}$ can be taken to be any distribution in $\mathcal{P}(\mathcal{Y})$. In particular, $U_{\alpha}$ is additive for $\alpha\in(0,1)\cup(1,\infty)$.
@@
Let us consider the case $\alpha\in(0,1)$ and rewrite
$\hskip 0.0ptU_{\alpha}(X;Y)$ $=\min_{Q_{Y}}\frac{1}{\alpha-1}\log\sum_{\begin{subarray}{c}x\in\mathcal{X}\\ y\in\mathcal{Y}\end{subarray}}P_{X}^{\alpha}(x)Q_{Y}^{\alpha}(y)P_{XY}^{1-\alpha}(x,y)$ (42)
@@
$\hskip 0.0ptU_{\alpha}(X;Y)$ $=\frac{1}{\alpha-1}\log\max_{\begin{subarray}{c}f\geq 0,\\ \|f\|_{1/\alpha}=1\end{subarray}}\langle f,g\rangle=\frac{1}{\alpha-1}\log\|g\|_{\frac{1}{1-\alpha}}=-\log\left\|\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{XY}^{1-\alpha}(x,\,\cdot\,)\right\|_{\frac{1}{1-\alpha}}^{\frac{1}{1-\alpha}},$ (44)
@@
$\hskip 0.0ptf=\frac{\left(\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{XY}^{1-\alpha}(x,\,\cdot\,)\right)^{\frac{\alpha}{1-\alpha}}}{\left[\sum_{y\in\mathcal{Y}}\left(\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{XY}^{1-\alpha}(x,y)\right)^{\frac{1}{1-\alpha}}\right]^{\alpha}}\eqcolon\left(\ddot{P}_{Y}^{(\alpha)}\right)^{\alpha}.$ (45)
We can rewrite (44) further:
$\hskip 0.0ptU_{\alpha}(X;Y)$ $=-\log\sum_{y\in\mathcal{Y}}P_{Y}(y)\left(\sum_{x\in\mathcal{X}}P_{X}^{\alpha}(x)P_{X|Y}^{1-\alpha}(x|y)\right)^{\frac{1}{1-\alpha}}.$ (46)